OpenStudy (hockeychick23):

Probability theory predicts that there is 77.6% chance of a particular soccer player making 2 penalty shots in a row. If the soccer player taking 2 penalty shots is simulated 2500 times, in about how many simulations would you expect at least 1 missed shot?

OpenStudy (mathmate):

@hockeychick23 you there?

OpenStudy (hockeychick23):

yes

OpenStudy (mathmate):

If we let p=the probability of making one shot, then q=1-p is the probability of NOT making one shot.

OpenStudy (hockeychick23):

I did the problem but I just wasn't sure if I completed it correctly, i took .224 and multiplied it by 2500 to get 560. But I just wasn't sure if I took the correct steps

OpenStudy (hockeychick23):

and i got .224 by doing 1-.776

OpenStudy (mathmate):

.224 is the probability of not making TWO shots, which could be zero or one.

OpenStudy (hockeychick23):

oh ok, so would that mean i would have to multiply my answer by 1/2 since its two shots?

OpenStudy (mathmate):

When p is the probability of making ONE shot, the probability of making TWO shots would be p^2=0.776 by the multiplication law. Can you find p?

OpenStudy (hockeychick23):

yea, p=0.88090862182

OpenStudy (mathmate):

Good!

OpenStudy (mathmate):

So what is the probability of MISSING one shot?

OpenStudy (hockeychick23):

1-0.88090862182 which would be0.11909137817

OpenStudy (mathmate):

Great! You're doing well!

OpenStudy (mathmate):

Now, what is the probability of missing TWO shots?

OpenStudy (mathmate):

Recall the multiplication rule!

OpenStudy (mathmate):

Nooooo! The multiplication rule says that if q is the probability of missing one shot, the probability of missing TWO shots will be q^2!

OpenStudy (hockeychick23):

oh ok now i got .01418= .11909137817^2

OpenStudy (mathmate):

I've got to go. But if you interpret what (1-q^2) means, and work with n=2500, you would get the answer quite easily. I'll come back to check on your answer later in the evening.

OpenStudy (hockeychick23):

ok i got 1940 as my final answer

OpenStudy (hockeychick23):

but also considering theres a 77.6% chance they will make 2 in a row it makes much more sense that they would only miss 560 doesn't it?

OpenStudy (mathmate):

q is the probability that she will miss one shot. (q=0.11909137817819) and q^2 is the probability that she will miss both shots. Therefore 1-q^2 is the probability that she will make at least one shot. Since there are 2500 trials, each of which has probability 1-q^2 of making at least one shot, then approximately 2500(1-q^2) will be the expected number of trials that she gets at least one shot. If not clear, please post again. @hockeychick23

OpenStudy (hockeychick23):

Ye so I got q^2= 0.01418275635638067, then did 1-0.014182756356380670= .98581724364361933 then did.98581724364361933(2500) and got 2464.543109109048325

OpenStudy (hockeychick23):

but it seems way to big to be correct

OpenStudy (mathmate):

Indeed! I have mis-read the question, which asks for "missing at least one shot". This means that she did not get both shots, or 1-0.776=0.224, as you suggested earlier. This situation is represented by the shaded part of the figure below:  and the correct answer is 2500(1-0.776)=560 I apologize for having misread the question.